Examine the convergence of the following series:

i) (3×4)/52 + (5×6)/72 + (7×8)/92....


ii) 1 + 4x + 42x2 + 43x3 +....(x > 0)

1
Expert's answer
2021-06-07T16:05:05-0400
"\displaystyle\sum_{i=1}^{\infin}\dfrac{(2n+1)(2n+2)}{(2n+3)^2}=\dfrac{3\times4}{5^2}+\dfrac{5\times6}{7^2}+\dfrac{7\times8}{9^2}+..."

Use the Test for Divergence


"\lim\limits_{n\to \infin}a_n=\lim\limits_{n\to \infin}a_n\dfrac{(2n+1)(2n+2)}{(2n+3)^2}"

"=\lim\limits_{n\to \infin}a_n\dfrac{(2+\dfrac{1}{n})(2+\dfrac{2}{n})}{(2n+\dfrac{3}{n})^2}"

"=\dfrac{2(2)}{(2)^2}=1\not=0"

The given series diverges by the Test for Divergence.


(ii)


"\displaystyle\sum_{i=0}^{\infin}(4x)^n=1+4x+4^2x^2+4^3x^3+... (x>0)"

The geometric series


"\displaystyle\sum_{i=0}^{\infin}ar^n"

is convergent if "|r|<1."

If "|r|\geq1," the geometric series diverges.

We have "r=4x, x>0"

Then the given series converges for "0<x<\dfrac{1}{4}" and diverges for "x\geq \dfrac{1}{4}."